OpenStudy (anonymous):

Input the equation for the following line. Reduce all fractional answers to lowest terms. A line that passes midway between and is parallel to both x - y = 6 and x - y = -6. i got y=x-6

OpenStudy (anonymous):

wouldn't that by \[x-y=0\]?

OpenStudy (aroub):

y=-x

OpenStudy (anonymous):

i could be wrong because i just said half way between 6 and -6 is 0

OpenStudy (anonymous):

the 2 can be written y = x - 6 y = x + 6 y = x is the answer

OpenStudy (anonymous):

which is in fact y = x, not y = -x

OpenStudy (anonymous):

yep

OpenStudy (aroub):

yes sorry y=x